### v2

# Topological Symmetry Breaking in Antagonistic Dynamics

Topological Symmetry Breaking in Antagonistic Dynamics
Giulio Iannelli1,2    Pablo Villegas1,3 pablo.villegas@cref.it    Tommaso Gili4,1    Andrea Gabrielli1,5 1Enrico Fermi Research Center (CREF), Via Panisperna 89A, 00184, Rome, Italy 2Dipartimento di Fisica, Università di Roma “Tor Vergata”, 00133 Rome, Italy 3Instituto Carlos I de Física Teórica y Computacional, Universidad de Granada, Granada, Spain 4Networks Unit, IMT Scuola Alti Studi Lucca, Piazza San Francesco 15, 55100- Lucca, Italy. 5Dipartimento di Ingegneria Civile, Informatica e delle Tecnologie Aeronautiche, Università degli Studi “Roma Tre”, Via Vito Volterra 62, 00146, Rome, Italy
A dynamic concordia discors, a finely tuned equilibrium between opposing forces, is hypothesized to drive historical transformations 1. Similarly, a precise interplay of excitation and inhibition, the 80:20 ratio, is at the basis of the normal functionality of neural systems 2, 3, 4. In artificial neural networks, reinforcement learning allows for fine-tuning internal signed connections, optimizing adaptive responses to complex stimuli, and ensuring robust performance 5, 6, 7. At present, engineered structures of competing components are, however, largely unexplored, particularly because their emergent phases are closely linked with frustration mechanisms in the hosting network 8, 9. In this context, the spin glass theory has shown how an apparently uncontrollable non-ergodic chaotic behavior arises from the complex interplay of competing interactions and frustration among units 10, 11, leading to multiple metastable states preventing the system from exploring all accessible configurations over time 12. Here, we tackle the problem of disentangling topology and dynamics in systems with antagonistic interactions. We make use of the signed Laplacian operator to demonstrate how fundamental topological defects in lattices and networks percolate, shaping the geometrical arena and complex energy landscape of the system. This unveils novel, highly robust multistable phases and establishes deep connections with spin glasses when thermal noise is considered, providing a natural topological and algebraic description of their still-unknown set of pure states at zero temperature.
Competing interactions introduce intrinsic incompatibilities between fundamental dynamical interactions and the elementary scales of the underlying lattice geometry 13, shaping the collective behavior of the system. This gives rise to new fundamental phenomena and yields to intriguing effects, such as the formation of exotic states like spin ice 14, 15, spin liquids 16, and spin glasses that have tantalized physicists for the last 40 years 17, 8. The latter, the paradigm of disordered magnetic systems, are characterized by competing interactions among their constituent spins, leading to an intricate and unpredictable dynamical and statistical configurational behavior 11. Signed interactions induce frustration and play a crucial role in describing the complex dynamics of these systems.
The thorough study of spin glasses opened the door to one of the most elegant proposals in the field of modern theoretical physics: the replica trick and the replica symmetry breaking (RSB), which offers a comprehensive understanding of their non-ergodic thermodynamic properties 18, 19, 20 in the mean-field scenario. Solid efforts have been made to detect the nature of the spin glass phase 12, 10, 21. Therefore, significant progress in both analytical and numerical aspects has been made: the discovery of temperature and disorder chaos 22, stochastic stability 23, metastases 24, or the rigorous analysis of short-range spin glasses 25, 26. In a pioneering work, Amit et al. established the rigorous connection between neural networks and Ising spin glasses, highlighting that both systems share an energy landscape with numerous metastable states and valleys that enable the delocalized storage of patterns 24.
However, several essential questions remain to be solved. Little is known about the number of pure ground states or the nature of broken symmetry in short-range spin glasses 27. Moreover, determining the ground states of Ising spin glasses in three or more dimensions is an NP-hard problem 28, which is closely related to many other hard combinatorial optimization problems 29. Similarly, disentangling new ensembles of correlated disorder 8, that is, predicting spin glass states in networks, represents an open challenge that remains crucial. In arbitrary architectures, in turn, it has never been fully understood the interplay between topology and signed interactions nor modeled such a delicate balance to generate multiple local states. Specifically, understanding the impact of competing units, i.e., elementary defects, in regular lattices and complex networks is key to properly controlling quantum error-correcting codes 30 and developing innovative brain-inspired devices, or even entirely new classes of structured artificial intelligence models that allow for more efficient and richer scenarios of machine-learning algorithms.
Canonical formulation of signed architectures
The multiscale structure of any homogeneous or heterogeneous signed network or lattice with antagonistic interactions can be analyzed using its corresponding Laplacian density matrix 31, 32, 33,
$\hat{\rho}(t)=\frac{e^{-t\bar{L}}}{Z(t)}$ (1)
where $Z(t)=\operatorname{Tr}[e^{-t\bar{L}}]$, $\bar{L}=\bar{D}-A$ represents the Signed Laplacian (SL) of the network 34, with $A$ being the signed adjacency matrix and $\bar{D}$ the diagonal matrix of elements $(\bar{D})_{ij}=\delta_{ij}\sum_{k}|A_{ik}|$. $\bar{L}$ can be derived from the usual reordering of the discrete Landau-Ginzburg Hamiltonian 35 and extends the Laplacian operator to signed architectures (see Methods).
Fig. LABEL:fig:a:lattc2sq_panel illustrates the SL in terms of spectral embedding arguments, as originally developed in 34 using the concept of antipodal proximity. Ordinary Laplacian embedding relies on the non-trivial solutions of the Laplace equation $L\bm{x}=0$ for placing vertices of a graph in Euclidean space. In this way, each vertex gets positioned at the weighted mean of its neighbors’ coordinates, with weights determined by the connecting edges. In contrast, when using $\bar{L}\bm{x}=0$ for embedding a signed graph, negatively connected neighbors contribute their coordinates with an opposite sign to the weighted sum determining the node placement. This makes nodes get closer to the antipodes of their oppositely tied neighbors. Moreover, since the SL has been formally proven to be a positive semidefinite operator (see Methods), we can rigorously implement thee recently introduced Laplacian framework 33, 32 to scan different spatial scales when an arbitrary architecture contains antagonistic or frustrated interactions. In particular, the heat capacity of a signed graph
$C=-\frac{\mathrm{d}S}{\mathrm{d}(\log\tau)},$ (2)
defined as the derivative of the graph entropy 31, 32, 33 $S(t)=\operatorname{Tr}[\hat{\rho}(t)\ln\hat{\rho}(t)]$ (see Methods and Supplementary Information (SI)) as a function of the (logarithm of the) diffusion time $\tau$ is, therefore, a detector of transition points that reveals the intrinsic characteristic diffusion scales of the system. Indeed, each pronounced peak of $C$ reveals a strong deceleration of the information diffusion, separating regions with strong communicability from the rest of the network where the diffusion slows down 32, 33.
Fig. LABEL:fig:b:lattc2sq_panel shows a sketch of a 2D lattice with $M$ edges, where a random fraction of negative edges, denoted by $p=M^{-}/M$, is introduced, being $M^{-}$ the total number of negative edges. Fig. LABEL:fig:c:lattc2sq_panel shows the average heat capacity $\overline{C}$ for different values of $p$ in 2D squared lattices. In particular, we observe that the peak at short diffusion times disappears as the number of negative links increases, disrupting the lattice translational invariance. Note that the point where $\overline{C}$ exhibits a peak at short times formally corresponds to the so-called ultraviolet cutoff $\Lambda$, i.e., the lattice spacing 33, 36, 37. The existence of this peak, together with the subsequent plateau, is closely related to the existence of translational invariance in the network38, 37, which is lost at some specific value $p_{\rm c}$ when negative links gradually replace the positive ones. The probability distribution of the first maximum of $C$ for different realizations of 2D lattices with variable levels of frustration is reported in Fig. LABEL:fig:d:lattc2sq_panel, exhibiting bistable behavior and a Maxwell-like point for $p_{\rm c}\approx 0.1$. This symmetry breaking in the very topology plays, as discussed below, a key role in determining, altering, and disrupting dynamic phases and phase transitions.
Figure 1: Signed Laplacian. (a) The antipodal proximity concept in the 2D spectral embedding with the SL 34. The position $\bm{x}_{i}$ of the $i^{\mathrm{th}}$ node is computed as the average position of the neighbors $\bm{x}_{j}$ weighted by the adjacency index $A_{ij}$. When $\operatorname{sign}(A_{ij})=-1$, the position $\bm{x}_{i}$ is updated in the direction opposite to $\bm{x}_{j}$. Black arrows correspond to coordinate axes. (b) Sketch of a 2D square lattice with random positive (blue links) and negative (red links) interactions. (c) Heat capacity, $C$, versus the temporal resolution parameter of the system, $\tau$, for different values of $p$ averaged over multiple disorder realizations (see legend). (d) Probability distribution, $-\log[P(\tau^{(0)}_{\max})]$, for the position, $\tau^{(0)}_{\max}$, of the first detected maximum of $C$ at varying $p$. Its shape changes from a single minimum for low $p$ to a bistable configuration at $p_{\rm c}$.
Figure 2: Percolation of topological defects in regular structures. (a) Partitioning of a square 2D lattice as indicated by the positive and negative signs of $|\lambda_{0}\rangle.$ A set of fractal clusters emerges beyond the critical value, $p_{\rm c}$. (b) The lower part of each figure shows the partitioning of a 3D lattice as indicated by the positive and negative signs of $|\lambda_{0}\rangle$ (blue and red nodes, respectively). The upper part of the figure enhances only the positive ones. (c) Order parameter, $P_{\infty}$ versus the fraction of negative links, $p$ for a squared 2D lattice of variable size (see legend). Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, $p$. Note that $\chi$ diverges as the size of the system increases. Black dashed lines represent the expected value for the spin-glass transition at $T=0$. (d) Cluster size distribution for a 2D square lattice of different system sizes (see legend). The black dashed line represents the estimated Fisher exponent, $\tau\approx 1.8$. (e) Order parameter, $P_{\infty}$ versus the fraction of negative links, $p$, for a 3D cubic lattice of variable size (see legend). Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, $p$. Note that $\chi$ diverges as the size of the system increases. Black dashed lines represent the expected value for the spin-glass transition at $T=0$. (f) Cluster size distribution for a cubic lattice of different system sizes (see legend). The black dashed line represents the estimated Fisher exponent, $\tau=2.18$. (g) Sketch of topological defects in a squared lattice: single links (S), Z defects representing the elementary cell, and X defects representing “anti-nodes”. (h) $P_{\infty}$ versus the fraction of Z defects, $p$ for a squared 2D lattice of variable size (see legend). Inset: Variance of the order parameter, $\chi$, as a function of the fraction of Z defects, $p$. Black dashed vertical line corresponds to $p_{\rm c}=0.1$. (i) $P_{\infty}$ versus the fraction of X defects, $p$ for a squared 2D lattice of variable size (see legend). Inset: Variance of the order parameter, $\chi$, as a function of the fraction of X defects, $p$. Black dashed vertical line corresponds to $p_{\rm c}=0.1$.
Percolation of topological defects
We still lack any indicator of a structural phase transition, as the observed fraction of negative links does not have an immediate interpretation in terms of site or bond percolation problems 39. A thorough analysis of the SL by exploring the parallel with the heat-like equation $\partial_{t}\psi=-\bar{L}\psi$ in the continuum limit approximation, leads to the expression $\bar{L}\equiv-\operatorname{\nabla^{2}}+B$, where $B$ is an arbitrary operator. This supports the analysis of any signed network as a perturbation to the usual network Laplacian, being formally analogous to the case of point-like particles diffusing in a uniform medium and being absorbed by a single sink (see Methods). In particular, it can be easily demonstrated that a negative link generates an “effective” influence area, affecting the surrounding region. It comes natural to speculate that the inclusion of single isolated negative links can be interpreted as a continuum percolation problem, where the influence area will depend on the specific underlying topology. Although this illustration does not resolve our original problem, it offers a valuable physical interpretation of the phenomena that occur as antagonistic interactions are progressively introduced to the system.
Inspired by the fundamental stabilizers proposed by Kitaev in the realm of topological computation (essential for error detection and correction in the surface code model 30), we propose and analyze the three fundamental topological defects: X defects, that can be mapped to nodes with all negative links, symbolizing disrupted interactions, Z defects, that correspond to frustrated plaquettes, or elementary system cells (see Fig. LABEL:fig:g:percolation_panel), and the already discussed single defects (S) corresponding to negative links. From now on, $p$ corresponds to the relative number of defects included in the system.
We propose to use the components of the eigenvector ($|\lambda_{0}\rangle$) corresponding to the smallest eigenvalue $\lambda_{0}$ of $\bar{L}$ to define a universal order parameter for the emergence of a topological phase transition. Specifically, we take the signs of $|\lambda_{0}\rangle$ to identify independent mesoscopic clusters, with the most abundant elements assigned a positive sign. Note that in the limiting case $p=0$, $|\lambda_{0}\rangle$ is a uniform vector, but as $p$ increases, it develops a mix of positive and negative components, capturing how the system fragments into separate clusters. This transition naturally connects to percolation theory, offering a general framework to analyze it.
In particular, $|\lambda_{0}\rangle$ captures how the system fragments when global translational invariance –and with it, the unique ultraviolet cutoff, $\Lambda$– is lost. As a result, the ordered (ferromagnetic) phase spanning the entire system, represented by a uniform vector, is no longer a physically viable solution. Fig. LABEL:fig:a:percolation_panel illustrates different system configurations at criticality, above, and below the expected percolation critical threshold for 2D square lattices, while Fig. LABEL:fig:b:percolation_panel presents the same for 3D cubic lattices when S defects are introduced. We must make just one final technical remark: since in regular lattices with $p=0$, $\lambda=k^{2}$, small eigenvalues correspond to large spatial scales in $k-$ space. Thus, numerical precision sets a resolution limit on the maximum system size that can be reliably analyzed for a given topology. This constraint arises from double-precision arithmetic, as discussed in detail in SI.
Fig. LABEL:fig:c:percolation_panel shows the percolation phase transition for a 2D square lattice where a double-peak percolation transition emerges. There, the fraction of points belonging to the largest cluster, $P_{\infty}=n_{\infty}/N$, is the order parameter, with $N$ the total system size and $n_{\infty}$ the number of sites belonging to the percolating cluster. The intrinsic fluctuations, $\chi\equiv\sqrt{N}\sigma_{P_{\infty}}$, diverge at $p_{\rm c}$ as expected for usual second-order phase transitions 40. As reported in the inset of Fig. LABEL:fig:c:percolation_panel, the first peak in the susceptibility, located at $p_{\rm c}\sim 0.10(1)$, is fully compatible with the expected threshold for frustrated plaquettes and/or the emergence of the spin glass phase at zero temperature 41. Further examples considering triangular and hexagonal lattices are reported in SI, predicting the threshold where spin glass phases emerge 41, 42 (we will discuss this important point below). Moreover, Fig. LABEL:fig:e:percolation_panel shows the percolation phase transition for a 3D cubic lattice, where a double peak in the system susceptibility emerges, with $p_{\rm c}=0.22(1)$ 43 (4D analysis are reported in SI). The cluster-size distribution (see Fig. LABEL:fig:d:percolation_panel and Fig. LABEL:fig:f:percolation_panel) confirms that the phase transition effectively belongs to the ordinary percolation universality class, as $P(S)\sim S^{-\tau}$, with $\tau\approx 2.18(1)$ for the 3D case. The case of 2D lattices is more subtle, as we observe an exponent $\tau\approx 1.75(1)$, still compatible with previous results in small-to-mid systems in continuum percolation problems, where logarithmic corrections may play a key role in the convergence of the cluster size distribution for low dimensions 44.
The double peak in the susceptibility represents a delicate point that must be discussed in terms of the geometrical properties of percolation in Euclidean spaces. Specifically, in three dimensions, two spanning clusters can geometrically coexist due to the natural emergence of lower-dimensional percolation structures. This is a natural consequence of the Alexander-Orbach conjecture 45: as critical percolation clusters present a universal spectral dimension $d_{S}=\nicefrac{{4}}{{3}}$ they can only coexist in dimension $d\geq 3$. In fact, we observe how the double-peak structure represents a finite-size effect in 2D lattices, which disappears in the infinite-size limit. Instead, in 3D, this phenomenon survives in the thermodynamic limit (a fact that is fully compatible with recent observations of the coexistence of spin glass and ferromagnetic order in spin glass 3D systems 46).
We report in Fig. LABEL:fig:h:percolation_panel and Fig. LABEL:fig:i:percolation_panel the phase transition when, respectively, Z or X defects are included in a 2D lattice. In particular, for both cases, the percolation threshold is equivalent to those of S defects, as they exhibit a percolation phase transition for the same values of $p_{\rm c}$ (see also SI for further analyses on different lattices). The case of X defects generates an entirely new situation. In fact, in the squared 2D lattice, the double-peak structure does not vanish in the thermodynamic limit, as the second peak is directly related to the site-percolation problem. Note that X defects in the low-density regime are not expected to produce any closed loop, leading to no local frustration of interactions (the system is, then, effectively balanced, i.e., the fundamental eigenvalue is equal to zero). This leads to new fundamental dynamic implications that are explored below when X defects are considered together with a dynamic process.
Figure 3: Ising dynamics. (a) Qualitative $T-p$ phase diagram for the 2D squared lattice with S defects. The lower part shows the energy levels corresponding to the first 100 eigenstates of $\bar{L}$ (see colorbar), for two values of $p$, respectively, in the ferromagnetic phase and the spin-glass one. The lowest eigenstate of $\bar{L}$ has been marked in red. Note that, inside the spin-glass phase, all eigenstates come closer in energy, and their ordering is lost at the ’onset of degeneracy’ at $p_{\rm c}$. (b) Qualitative $T-p$ phase diagram for X defects computed with the SL eigenstates through numerical simulations in a 2D square lattice of size $N=64^{2}$. The breakup of the ferroelectric phase translates into the emergence of a metastable one where the SL eigenvectors now define multiple degenerate stable states characterized by local order up to the paralectric/ferroelectric phase transition. At a certain $p$, stable minima becomes unstable, leading to a spin-glass phase. The lower part shows the energy levels corresponding to the first 100 eigenstates. In this case, the energy degeneracy is not observed before $p=0.4$, where the spin-glass phase emerges. (c) Annealing and sudden quenching initializing the system in a low eigenmode $|\lambda_{2}\rangle$ spin configuration on a 2D squared lattice with $p=0.25$ fraction of X defects. We heat the system up to $T\approx 1.5$, i.e., smaller but closer to the critical Ising temperature. After relaxation, the system is suddenly quenched, observing that, up to border effects, it does not escape from the local minima associated to $\operatorname{sign}(|\lambda_{2}\rangle)$. (d) Same as the previous case but using S defects with the lowest SL eigenmode at the $p_{\rm c}$. In this case, the system is locally stable but starts the attractor surfing at much lower temperatures. (e) Local magnetization order parameter, $m^{\rm GC}$, as a function of the temperature and the fraction of X defects for $N=64\times 64$. (f)- (g) Spin overlap, $\langle\lambda_{i}\,|\,\mathopen{}\sigma_{\infty}\rangle$, as a function of the temperature for different lattice sizes for (f) X defects, using $|\lambda_{2}\rangle$, and (g) S defects, using $|\lambda_{2}\rangle$. The critical temperature is indicated in the insets as a vertical dashed line. Note how $T_{\rm c}$ is significantly greater for the case of X defects, while, in contrast, the initial configuration loses its stability at much lower temperatures for S defects. All curves have been averaged over $n_{\rm avg}=1000$ independent realizations.
Exploring the dynamical arena of antagonistic non-ergodic systems
In frustrated systems, the emergence of non-ergodic dynamics manifests in the existence of multiple stationary states with disjoint basins of attraction 9. Hence, predicting stationary states in such systems at low temperatures represents a significant challenge due to the intrinsic geometrical complexity of the problem. We state the following conjecture: for a vanishing thermal noise, i.e., at zero temperature, these effects are fully described by the topological symmetry-breaking mechanism reported in the previous section. Their emergence will thus be strictly linked to the percolating clusters defining the structural shape of the system. Specifically, we have analyzed the dynamical implications of the different topological defects simulating the Edwards-Anderson (EA) Hamiltonian (see Methods),
$\mathscr{H}=-\sum_{(i,j)}J_{ij}\sigma_{i}\sigma_{j},$ (3)
where $\sigma_{i}$ is the spin in the $i^{\rm th}$ site, with $J_{ij}=JA_{ij}$, being $A$ the signed adjacency matrix, and $J$ the global coupling strength.
Fig. 3(a) and (b) show the temperature phase diagram of the system as a function of the fraction of negative links, $p$, for S and X defects on 2D square lattices, respectively. In particular, S defects correspond to the well-known Random Bond Ising Model (RBIM) or ±$\pm J$-model 47, 48 (see Methods). Results are achieved through extensive Monte Carlo simulations of spin-spin dynamics, running Eq. (3) on top of the signed structure. The inspection of the phase diagram confirms the well-known collective dynamical regimes for S defects: ferromagnetic and spin glass phases. Now, we characterize the eigenstates of the system using the corresponding signed eigenvectors of $\bar{L}$ to define topologically-based up and down spin islands. Then, we calculate the energy of the system at $T=0$ through the EA Hamiltonian as defined in Eq. (3). This confirms the degeneracy of energetic states at the critical fraction of negative links, $p_{\rm c}$, for both S and X defects. Fig. 3(a) and (b) report how the ground state energy gap disappears, leading to the degeneracy and, finally, inversion of the energetic levels of the system. Note that, for X defects, the diagram shows, however, distinct bifurcation lines separating new phases that are still unknown to the best of our knowledge. In particular, for low temperatures, increasing the fraction $p$, a first transition separates the ferromagnetic phase from a metastable one, where the system can get trapped into local minima up to the Ising critical temperature, $T_{\rm c}$ (see Fig. 3(c)). A second transition at higher $p$ leads to the spin glass phase, where the resilient and locally stable spin islands become unstable when thermal noise is applied. Notably, this new intermediate phase between the ferromagnetic and spin glass phases, the MM phase, exhibits multiple coexisting local islands of mesoscopic magnetic order. Note that, unlike the chaotic attractor “surfing” seen in spin glass phases at low temperatures, the MM phase allows access to highly diverse noise-resilient functional configurations.
Note that, for S defects, the usually reported intriguing re-entrance of the ferroelectric phase to slightly higher values of $p$ when the thermal noise increases 49 can be ascribed to the breaking of small frustrated clusters due to noise effects. Specifically, thermal noise induces random flips with a single-spin interaction radius that is explicitly smaller than the size of the clustered defects. This effect first breaks small spin-clusters, reducing domain wall effects and making more robust the giant ferroelectric cluster.
Fig. 3(c) and (d) illustrate the stability of different eigenstates of $\bar{L}$ as the system is heated up to different temperatures. We emphasize that, for X defects, the system can be heated up to the Ising critical temperature, where it is reset by noise, and allow for choosing other possible local minima. The same occurs at a much lower temperature, $T_{\rm SG}(N)\ll T_{\rm c}^{\rm Ising}$, for S defects. In the latter case, typical of spin glass behavior, the minima are locally stable, but the system can escape under low noise levels, exploring the phase space (see SI). We hypothesize that the spin-glass tree of states 9 is expected to be a superposition of the $N$ eigenvectors of the SL, or pure states, that represent the original set of local minima.
It comes natural to introduce topologically-based local order parameters as a function of $T$ that effectively capture the emergent mesoscopic order when spontaneous symmetry breaking forms magnetized spin islands (see Methods). Fig. 3(e) shows the magnetization of the giant cluster of $|\lambda_{2}\rangle$ as initial state, as a function of $p$ and $T$ for the case of X defects. Hence, Fig. 3(f) and (g) show the spin overlap $\langle\lambda_{i}\,|\,\mathopen{}\sigma_{\infty}\rangle$ between the initial state and the long-time dynamics when the system has thermalized (see Methods). In particular, for the case of X defects, this evidences how when the system is initialized at $|\lambda_{2}\rangle$, it undergoes a second-order phase transition with a critical temperature a little below the Ising critical one, $T_{\rm c}\simeq 2.27$. This provides direct evidence of the stability of the corresponding eigenstate or initial configuration of the system. By contrast, Fig. 3(g) shows the narrow temperature range in which the selected starting eigenstates for S defects lose their stability, causing the system to start surfing across other attraction basis, i.e., pure states, or combinations thereof.
I Topological defects in heterogeneous architectures
To establish whether the SL is able to predict a percolative phase transition in heterogeneous systems, we have analyzed different dilute models and network structures by introducing a random fraction of S defects. We highlight that this approach generalizes the detection of frustrated dynamical phases to complex networks and generic disordered systems. Specifically, the results for Erdős-Rényi (ER) and Small World (SW) networks are reported in Fig. 4(a) and (b). Additionally, we have analyzed a particularly complex case with significant physical relevance: the 3D diluted cubic lattice, which has originally inspired dilute Ising models 50. Now, a fraction $f$ of the existent links are removed, remaining below the edge-percolation threshold to ensure the system is connected. We then assign positive and negative weights to the remaining interactions, randomly drawn from a Gaussian distribution with mean $\mu$ and variance $\sigma$. As illustrated in Fig. 4(c), we also predict the emergence of the frustrated phase as a function of $\sigma$. Finally, Fig. 4(d) shows the case of an ER network with X defects reconstructed with the local Ising order parameter.
Figure 4: Heterogeneous architectures. Order parameter, $P_{\infty}$ versus the fraction of negative links, $p$ for a: (a) ER network with $\langle\kappa\rangle=10$, (b) Small World network from a 2D lattice (rewiring probability $p_{\rm rew}=0.05$) and (c) a diluted 3D lattice ($p_{\rm dil}=0.25$) with normally weighted edges $p\sim\mathcal{N}(1,\sigma)$. Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, $p$. (d) $T-p$ Phase diagram of the ER network with X defects reconstructed with the local Ising order parameter $m^{\rm GC}_{\infty}$ measured on $\operatorname{sign}(|\lambda_{2}\rangle)$.
II Discussion and conclusions
We have solved a decades-long controversy surrounding the nature of spin glasses. We have explicitly shown how non-ergodic phases originated by competing interactions, such as the emergence of the spin glass phase at zero temperature, lie in a pure topological symmetry-breaking mechanism that goes beyond the considered dynamics. This mechanism exhibits a clear algebraic signature in the emergence of degeneracy of the ground state of the SL operator. Topological defects generate natural barriers in the discrete lattice that lead to the fragmentation of the system in mesoscopic regions already at low values of $p$. Naturally, such a mechanism strongly depends on the discrete nature of the underlying lattice, leading to different percolation thresholds even if the spatial dimension of the system remains unchanged. Hence, the SL framework offers a natural approach for detecting spin glass phases in any homogeneous or heterogeneous structure. This has profound implications for considering future dynamical Renormalization Group (RG) analysis as it directly acts on the Green function of the system by adding only a new term correcting $\operatorname{\nabla^{2}}$, which hampers the topological properties of the embedding space on top of which any field-theoretical approach is based.
It is precisely the local nature of particular defects that make different non-ergodic systems strongly unstable. For example, in the case of S defects that lack local stability, frustrated loops of arbitrary size can be generated by thermal noise, leading to large avalanches that cause the system to escape from a stable minimum, allowing it to explore large parts of the available space. In other words, frustrated loops are expected to induce noise effects, leading the system to a sempiternal out-of-equilibrium condition that permits it to make attractor surfing easier. This last case corresponds to the Z and S topological defects, making spin glasses elusive when searching for stable configurations. A different picture can be drawn for X defects: the system shows multiple minima now, but high energy barriers allow the system to remain trapped in each one, making it predictable. However, we want to emphasize that all defects generally lead systems to ergodicity breaking, where a single stable state no longer exists.
The percolative analysis presented here also permits an immediate interpretation of the long-standing debate of the upper critical dimension of spin glass systems 9. That is, on determining the dimensionality $d_{u}$ above which the critical behavior of spin glasses is described by mean-field theory. Suppose one assumes that the dimension of the percolative giant cluster determines the effective dimension where the spin system operates. In that case, $d=6$ is the upper dimension for the isotropic percolation universality class, which presents a fractal dimension $d_{f}=4$, which is precisely the upper dimension for the Ising universality class. Hence, specific spatial dimensions $d=4$ and $d=5$ are known to exhibit a fractal cluster with $d_{f}\approx 3.04$ and $d_{f}\approx 3.52$, respectively 51, making invalid the mean-field solution for Ising dynamics when competing interactions are considered.
Finally, our results significantly progress in defining natural stable states, determined by the SL operator, where frustrated systems can exhibit stable minima. Moreover, we emphasize that this step helps to reduce an NP problem –solved by hit-or-miss methods– to a combinatorial one of assembling system eigenstates as fundamental building blocks.
Our framework also opens a route to extend this approach to learning algorithms, helping to develop a common mathematical framework that better characterizes non-ergodic phases that facilitate engineered computation depending on specific microscopic details.
III Acknowledgments
P.V. acknowledges the Spanish Ministry of Research and Innovation and Agencia Estatal de Investigación (AEI), MICIN/AEI/10.13039/501100011033, for financial support through Project PID2023-149174NB-I00, funded also by European Regional Development Funds, as well as Ref. PID2020-113681GB-I00. We also thank G. Cimini, M.A. Muñoz, P. Caputo, and C. Bordenave for their useful suggestions and comments.
IV Methods
The discrete signed Laplacian.
The traditional definition of the Laplacian operator, represented by $L=D-A$, presents serious issues when dealing with signed networks as it is no longer semi-definite positive. This is due to the fact that diagonal terms may take null or even negative values and $D^{-1/2}$ and, therefore, $L_{\rm RW}$ is no longer defined. To tackle this delicate point, a recent proposal has mathematically introduced the signed version of the Laplacian operator. It is grounded in sound physical principles, being capable of accurately capturing the complex and intricate diffusion dynamics of any network 34. In particular, the SL of a signed architecture can be defined as,
$\bar{L}=\bar{D}-A,$ (4)
where $\bar{D}$ represents the unsigned degree matrix, i.e., the diagonal matrix given by $\bar{D}_{ii}=\sum\nolimits_{j}|A_{ij}|$, with $A$ the signed adjacency matrix.
From Eq. (4), it follows that the SL can be rewritten as the usual Laplacian $\bar{L}$ of the unsigned version of the network minus two times the adjacency matrix of the negative links $A^{-}$, namely,
$\bar{L}=L-2A^{-},$ (5)
which, in the continuum approximation, can be seen as
$\bar{L}\equiv-\operatorname{\nabla^{2}}+B.$ (6)
where $B$ represents a generic operator affecting the usual second derivative term, tantamount to sink effects.
Properties of the signed Laplacian.
The SL operator preserves relevant properties, namely:
  1. (i)
    The SL matrix $\bar{L}$ is positive-semidefinite for any graph.
  2. (ii)
    $\bar{L}$ becomes strictly positive-definite for the unbalanced graphs (that is, for connected unbalanced networks, there is only one eigenvalue strictly equal to zero, $\lambda_{\min}=0$, in the total absence of negative loops).
  3. (iii)
    $\lambda_{\min}$ measures the level of balance (frustration) of the network 34.
SL as diffusion dynamics in the presence of sinks.
In particular, Eq. (6) can be reduced, when considering only one negative link, to the problem of point-like particles diffusing in a uniform medium and being absorbed by a single spherical sink 52 (see further details in SI), with $m=m(\bm{x},t)$ the magnetization field,
$\frac{\partial m}{\partial t}=D_{0}\operatorname{\nabla^{2}}m-\lambda(\bm{x})m+s(\bm{x},t),$ (7)
where $D_{0}$ stands for the diffusion coefficient, $\lambda(\bm{x})$ is the local absorption coefficient and $s(\bm{x},t)$ the source density.
Field-theoretical implications.
It is well-known that the Green function for the Gaussian model follows the expression,
$G(\textbf{k},\omega)=[-i\omega+D_{0}k^{2}]^{-1}$ (8)
Through the extension of Eq. (7) to many sinks 52 one can find the following Green function,
$G(\textbf{k},\omega)=[-i\omega+D_{0}k^{2}+nt(\textbf{k},\textbf{k},\omega)]^{-1}$ (9)
where the last term represents the interaction with sinks in the system. This opens the door to the rigorous formulation of a field-theoretical framework for spin glasses and other frustrated dynamics as, in principle, the signed interactions add an extra term to the Green function, which is the basis of diagrammatic expansions of the correlation function.
Canonical formulation of signed interactions.
Once we can consider a semidefinite operator for signed networks, $\bar{L}$, it is possible, in terms of the network propagator, $\hat{K}=e^{-\tau\bar{L}}$, to generalize the ensemble of accessible information diffusion states 31, 33, 32, 53, namely,
$\hat{\rho}(\tau)=\frac{e^{-\tau\bar{L}}}{Z},$ (10)
where $\rho(\tau)$ is tantamount to the canonical density operator in statistical physics (or to the functional over fields configurations) 35, 54, and $Z(\tau)=\sum_{i}^{N}e^{-\tau\lambda_{i}}$, being $\lambda_{i}$ the set of system’ eigenvalues. It is possible to, therefore, define the network entropy 31 through the relation
$S[\hat{\rho}(\tau)]=-\operatorname{Tr}\left[\hat{\rho}(\tau)\log\hat{\rho}(\tau)\right]=\tau\langle\lambda\rangle_{\tau}+\log Z(\tau)$ (11)
being $\langle\hat{O}\rangle_{\tau}=\operatorname{Tr}[\hat{\rho}(\tau)\hat{O}]$. Immediately, it is possible to define the heat capacity of any signed network as 32,
$C=-\frac{\mathrm{d}S}{\mathrm{d}(\log\tau)}=-\tau^{2}\frac{\mathrm{d}\langle\lambda\rangle_{\tau}}{\mathrm{d}\tau}$ (12)
Eq. (12) describes the heat capacity of the network, which can detect transition points that correspond to the intrinsic diffusion scales characteristic of the network. When condition ${\mathrm{d}C}/{\mathrm{d}\tau}|_{\tau^{*}}=0$ is met, it defines $\tau^{*}$ and reveals pronounced peaks that reflect a significant deceleration in information diffusion. If $C$ also shows a well-defined plateau, it represents direct evidence of a scale-invariant architecture 33, 36, 38 (such as, for example, 2D lattices). In this case, a peak at short times, in turn, independent of system size, captures the ultraviolet cutoff, $\Lambda$ 37. This cutoff represents a minimum scale that prevents divergence in usual RG calculations 55 and indicates that the system exhibits some kind of translational invariance if the subsequent plateau is present 38.
Signed Ising dynamics.
The archetypal spin glass model is defined as the celebrated EA model 48, with the corresponding Hamiltonian
$\mathscr{H}=-\sum_{(i,j)}J_{ij}\sigma_{i}\sigma_{j},$ (13)
where $\sigma_{i}=\pm 1$ are Ising spins, $(i,j)$ are the nearest-neighbors’ bonds of the network and $J_{ij}=JA_{ij}$, with $A_{ij}$ the weighted and signed adjacency matrix. This represents the paradigmatic model for disordered magnets with short-range spin-spin interactions, including multiple negative loops.
Note that the previous equation can be safely rewritten as 35,
$\mathscr{H}=-J\sum_{(i,j)}A_{ij}\sigma_{i}\sigma_{j}+J\sum_{(i,j)}k_{i}\delta_{ij}\sigma_{i}\sigma_{j},$ (14)
where $k_{i}$ is the node connectivity. This automatically leads to the natural definition of the SL as written in Eq. (4).
We have set, for simplicity, $J=1$. For the case of S defects, the $ij^{\rm th}$ value is a random variable extracted from the probability density function $P(J_{ij})=(1-p)\delta(J_{ij}-J)+p\delta(J_{ij}+J)$, which correspond to the RBIM. Regarding X defects, the parameter $p$ denotes the probability that a site or node is selected for flipping, which entails that all edges incident on that node are assigned a negative weight. In contrast, for Z defects, the parameter $p$ represents the fraction of unit cells, i.e., closed loops within the lattice in which every edge is flipped.
The global normalized magnetization of the system is naturally defined as,
$m=\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}$ (15)
where $N$ corresponds to the system size and $s_{i}$ is the state of the Ising spins at site $i$.
Instead, we can measure the local order of the system through different observables when using some selected SL eigenstates, $|\lambda_{i}\rangle$ as initial states. To perform this, the system is prepared in a selected ’pure’ eigenstate, where negative and positive signs are associated to up and down states, respectively.
One possibility corresponds to the so-called spin overlap, i.e. the superposition between the initial state and the dynamical evolution of the system at each time $t$, as follows,
$\langle\lambda_{i}\,|\,\mathopen{}\bm{\sigma}(t)\rangle=\frac{1}{N}\sum_{k=1}^{N}\delta(\operatorname{sign}(\lambda_{i}^{(k)})-\sigma_{k}(t))$ (16)
where $\lambda_{i}^{(k)}$ is the k-th component of the i-th eigenmode $|\lambda_{i}\rangle$, with $\bm{\sigma}$ the vector of magnetizations at time $t$.
Alternatively, one can measure the magnetization of the giant cluster of the system in its initial state. Then, the local order parameter runs over the $n_{\infty}$ spins of the giant component of this specific state as indicated by $|\lambda_{i}\rangle$,
$m^{(i)}_{\infty}(t)=\frac{1}{n^{(i)}_{\infty}}\stackrel{{\scriptstyle[}}{{j}}=1]{n^{(i)}_{\infty}}{\sum}\sigma_{j}(t)$ (17)
RBIM vs field theory at $T\to 0$.
In order to tighten the relation between the fundamental state of the SL and the ground state of the Hamiltonian Eq. (3) with the random interaction matrix $J_{ij}$, it is convenient to introduce the field theoretical formulation of the Ising model given by the Hubbard-Stratonovich transformation 56 for which
$Z(J)$ $=\sum_{\bm{\sigma}}e^{-\beta\mathscr{H}[\bm{\sigma}]}$ (18)
$\propto\int\mathcal{D}\phi\,\exp\bigg{\{}-\beta\sum_{(i,j)}J_{ij}\phi_{i}\phi_{j}$
$\quad+\sum_{i}\ln\Big{[}\cosh\Big{(}2\beta\sum_{j}J_{ij}\phi_{j}\Big{)}\Big{]}\bigg{\}},$
where $\beta=1/T$ and $\mathcal{D}\phi=\prod_{i}d\phi_{i}$. For $\beta\to+\infty$ we can apply the saddle-point method leading to the self-consistent equations for the ground state:
$\sum_{i}J_{ki}\left(\phi_{i}-\operatorname{sign}\left[\sum_{l}A_{il}\phi_{l}\right]\right)=0,$
for $k=1,2,....,N$, where $A_{ij}=J_{ij}/J$ is the adjacency matrix of the network of interactions. The self-consistent solution of the above equation is
$\phi_{i}=\operatorname{sign}\left[\sum_{l}A_{il}\phi_{l}\right].$ (19)
Note that, as expected, for the case $p=0$ this gives two specular ferromagnetic solutions either $\phi_{i}=+1$ or $\phi=-1$ for all $i$. The symmetry can be broken either by a small and uniform external magnetic field or by suitable boundary conditions.
At small $p\neq 0$ clearly two ferromagnetic ground states related by the same global up-down symmetry are still present, but they are no more homogeneous due to the inhomogeneous pair interactions. When instead $p$ becomes large enough, i.e. at the $T=0$ critical value $p_{c}$, the ferromagnetic ground state is no more unique, but many different and degenerate ground states are solutions of Eq. (19) and a spin glass transition appears.
Let us now see the relation between the solutions of Eq. (19) and the eigenvector(s) related to the least eigenvalue of $\bar{L}$. As aforementioned, it (they) satisfies the following equation
$\sum_{i=1}^{N}\bar{L}_{ij}\psi_{j}=\lambda_{0}\psi_{i}\,,$
which can be rewritten also as
$\sum_{i=1}^{N}A_{ij}\psi_{j}=(k_{i}-\lambda_{0})\psi_{i}\,,$ (20)
where $k_{i}=\sum_{j}|A_{ij}|$ is the coordination number of site $i$ ($k_{i}=2d$ for a $d-$dimensional square lattice). As shown in 57, $\lambda_{0}\leq\min_{i=1}^{N}k_{i}$. This implies that
$\operatorname{sign}[\psi_{i}]=\operatorname{sign}\left[\sum_{l}A_{il}\psi_{l}\right],$ (21)
for all connected networks, including regular lattices. Eq. (21) is formally the same of Eq. (19) with the fundamental difference that the solution of the former has to be found on the $N-$dimensional spherical surface of radius $\sqrt{N}$ while for the latter has to found on the vertices of the $N-$dimensional cube defined by $\phi_{i}=\pm 1$ for each $i=1,2,...,N$ (note however that in both cases $|\bm{\psi}|=|\bm{\phi}|=\sqrt{N}$. This implies that $\operatorname{sign}[\psi_{i}]$ can differ from $\operatorname{sign}[\phi_{i}]$ in some node in particular where there are large fluctuations in the values of $\psi$ between neighboring nodes. Anyway, being $\bm{\phi}$ and $\bm{\psi}$ characterized by the same normalization condition, we expect that the binarized vector $\bm{\psi}_{\rm bin}$ of components $\psi_{i}/|\psi_{i}|$ shows a similar behavior with respect the binary vector $\bm{\phi}$ on a suitable coarse-grained scale as large as the frustration walls.

### v3

@@
 In artificial neural networks, reinforcement learning allows for fine-tuning internal signed connections, optimizing adaptive responses to complex stimuli, and ensuring robust performance 5, 6, 7.
-At present, engineered structures of competing components are, however, largely unexplored, particularly because their emergent phases are closely linked with frustration mechanisms in the hosting network 8, 9.
+Engineered systems with antagonistic interactions remain comparatively unexplored, particularly because their emergent phases are closely linked with frustration mechanisms in the hosting network 8, 9.
 In this context, the spin glass theory has shown how an apparently uncontrollable non-ergodic chaotic behavior arises from the complex interplay of competing interactions and frustration among units 10, 11, leading to multiple metastable states preventing the system from exploring all accessible configurations over time 12.
-Here, we tackle the problem of disentangling topology and dynamics in systems with antagonistic interactions.
+Here, we show how topology constrains dynamics in systems with antagonistic interactions.
 We make use of the signed Laplacian operator to demonstrate how fundamental topological defects in lattices and networks percolate, shaping the geometrical arena and complex energy landscape of the system.
-This unveils novel, highly robust multistable phases and establishes deep connections with spin glasses when thermal noise is considered, providing a natural topological and algebraic description of their still-unknown set of pure states at zero temperature.
+This unveils novel, highly robust multistable phases and establishes deep connections with spin glasses when thermal noise is considered, providing a natural topological and algebraic description of emergent multistability and non-ergodicity in frustrated systems.
 Competing interactions introduce intrinsic incompatibilities between fundamental dynamical interactions and the elementary scales of the underlying lattice geometry 13, shaping the collective behavior of the system.
-This gives rise to new fundamental phenomena and yields to intriguing effects, such as the formation of exotic states like spin ice 14, 15, spin liquids 16, and spin glasses that have tantalized physicists for the last 40 years 17, 8.
+This gives rise to new fundamental phenomena and leads to intriguing effects, such as the formation of exotic states like spin ice 14, 15, spin liquids 16, and spin glasses that have tantalized physicists for the last 40 years 17, 8.
 The latter, the paradigm of disordered magnetic systems, are characterized by competing interactions among their constituent spins, leading to an intricate and unpredictable dynamical and statistical configurational behavior 11.
@@
 Solid efforts have been made to detect the nature of the spin glass phase 12, 10, 21.
-Therefore, significant progress in both analytical and numerical aspects has been made: the discovery of temperature and disorder chaos 22, stochastic stability 23, metastases 24, or the rigorous analysis of short-range spin glasses 25, 26.
+Therefore, significant progress in both analytical and numerical aspects has been made: the discovery of temperature and disorder chaos 22, stochastic stability 23, metastates 24, or the rigorous analysis of short-range spin glasses 25, 26.
 In a pioneering work, Amit et al.
 established the rigorous connection between neural networks and Ising spin glasses, highlighting that both systems share an energy landscape with numerous metastable states and valleys that enable the delocalized storage of patterns 24.
-However, several essential questions remain to be solved.
+However, several essential questions remain unanswered.
 Little is known about the number of pure ground states or the nature of broken symmetry in short-range spin glasses 27.
@@
 $\bar{L}$ can be derived from the usual reordering of the discrete Landau-Ginzburg Hamiltonian 35 and extends the Laplacian operator to signed architectures (see Methods).
-Fig.
-LABEL:fig:a:lattc2sq_panel illustrates the SL in terms of spectral embedding arguments, as originally developed in 34 using the concept of antipodal proximity.
+Figure LABEL:fig:a:lattc2sq_panel illustrates the SL in terms of spectral embedding arguments, as originally developed in 34 using the concept of antipodal proximity.
 Ordinary Laplacian embedding relies on the non-trivial solutions of the Laplace equation $L\bm{x}=0$ for placing vertices of a graph in Euclidean space.
@@
 This makes nodes get closer to the antipodes of their oppositely tied neighbors.
-Moreover, since the SL has been formally proven to be a positive semidefinite operator (see Methods), we can rigorously implement thee recently introduced Laplacian framework 33, 32 to scan different spatial scales when an arbitrary architecture contains antagonistic or frustrated interactions.
+Moreover, since the SL has been formally proven to be a positive semidefinite operator (see Methods), we can rigorously implement the recently introduced Laplacian framework 33, 32 to scan different spatial scales when an arbitrary architecture contains antagonistic or frustrated interactions.
 In particular, the heat capacity of a signed graph
@@
 Indeed, each pronounced peak of $C$ reveals a strong deceleration of the information diffusion, separating regions with strong communicability from the rest of the network where the diffusion slows down 32, 33.
+Figure LABEL:fig:b:lattc2sq_panel shows a sketch of a 2D lattice with $M$ edges, where a random fraction of negative edges, denoted by $p=M^{-}/M$, is introduced, being $M^{-}$ the total number of negative edges.
 Fig.
-LABEL:fig:b:lattc2sq_panel shows a sketch of a 2D lattice with $M$ edges, where a random fraction of negative edges, denoted by $p=M^{-}/M$, is introduced, being $M^{-}$ the total number of negative edges.
-Fig.
-LABEL:fig:c:lattc2sq_panel shows the average heat capacity $\overline{C}$ for different values of $p$ in 2D squared lattices.
+LABEL:fig:c:lattc2sq_panel shows the average heat capacity $\overline{C}$ for different values of $p$ in 2D square lattices.
 In particular, we observe that the peak at short diffusion times disappears as the number of negative links increases, disrupting the lattice translational invariance.
-Note that the point where $\overline{C}$ exhibits a peak at short times formally corresponds to the so-called ultraviolet cutoff $\Lambda$, i.e., the lattice spacing 33, 36, 37.
+Note that the point where $\overline{C}$ exhibits a peak at short times corresponds to the shortest resolvable topological interaction scale (by analogy with an ultraviolet cutoff), which in regular lattices coincides with the lattice spacing 33, 36, 37.
 The existence of this peak, together with the subsequent plateau, is closely related to the existence of translational invariance in the network38, 37, which is lost at some specific value $p_{\rm c}$ when negative links gradually replace the positive ones.
@@
 (c) Heat capacity, $C$, versus the temporal resolution parameter of the system, $\tau$, for different values of $p$ averaged over multiple disorder realizations (see legend).
-(d) Probability distribution, $-\log[P(\tau^{(0)}_{\max})]$, for the position, $\tau^{(0)}_{\max}$, of the first detected maximum of $C$ at varying $p$.
+(d) Probability distribution, $\log[P(\tau^{(0)}_{\max})$], for the position, $\tau^{(0)}_{\max}$, of the first detected maximum of $C$ at varying pp.
 Its shape changes from a single minimum for low $p$ to a bistable configuration at $p_{\rm c}$.
@@
 (c) Order parameter, $P_{\infty}$ versus the fraction of negative links, $p$ for a squared 2D lattice of variable size (see legend).
-Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, $p$.
+Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, pp.
 Note that $\chi$ diverges as the size of the system increases.
@@
 (e) Order parameter, $P_{\infty}$ versus the fraction of negative links, $p$, for a 3D cubic lattice of variable size (see legend).
-Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, $p$.
+Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, pp.
 Note that $\chi$ diverges as the size of the system increases.
@@
 (h) $P_{\infty}$ versus the fraction of Z defects, $p$ for a squared 2D lattice of variable size (see legend).
-Inset: Variance of the order parameter, $\chi$, as a function of the fraction of Z defects, $p$.
-Black dashed vertical line corresponds to $p_{\rm c}=0.1$.
+Inset: Variance of the order parameter, $\chi$, as a function of the fraction of Z defects, pp.
+The black dashed vertical line corresponds to $p_{\rm c}=0.1$.
 (i) $P_{\infty}$ versus the fraction of X defects, $p$ for a squared 2D lattice of variable size (see legend).
-Inset: Variance of the order parameter, $\chi$, as a function of the fraction of X defects, $p$.
-Black dashed vertical line corresponds to $p_{\rm c}=0.1$.
+Inset: Variance of the order parameter, $\chi$, as a function of the fraction of X defects, pp.
+The black dashed vertical line corresponds to $p_{\rm c}=0.1$.
+The heat capacity $C(\tau)$ serves as a spectral telescope: short times $\tau$ probe the high-energy (local) defects, while long times probe the low-energy (global) topological structure.
+Thus, anomalies in $C(\tau)$ allow us to detect the precise scale at which topological symmetry breaking occurs, providing a bridge to the structural analysis of defects that follows.
 Percolation of topological defects
 We still lack any indicator of a structural phase transition, as the observed fraction of negative links does not have an immediate interpretation in terms of site or bond percolation problems 39.
-A thorough analysis of the SL by exploring the parallel with the heat-like equation $\partial_{t}\psi=-\bar{L}\psi$ in the continuum limit approximation, leads to the expression $\bar{L}\equiv-\operatorname{\nabla^{2}}+B$, where $B$ is an arbitrary operator.
-This supports the analysis of any signed network as a perturbation to the usual network Laplacian, being formally analogous to the case of point-like particles diffusing in a uniform medium and being absorbed by a single sink (see Methods).
+A complementary physical interpretation comes from the heat-like equation $\partial_{t}\psi=-\bar{L}\psi$: near the defect-percolation threshold, the diverging correlation length motivates an effective coarse-grained description in which $\bar{L}$ behaves as a Laplacian perturbed by localized sink-like terms, i.e.
+$\bar{L}\approx-\operatorname{\nabla^{2}}+B$, where $B$ encodes the coarse-grained effect of antagonistic links (see Methods and SI).
 In particular, it can be easily demonstrated that a negative link generates an “effective” influence area, affecting the surrounding region.
-It comes natural to speculate that the inclusion of single isolated negative links can be interpreted as a continuum percolation problem, where the influence area will depend on the specific underlying topology.
+It comes natural to speculate that the inclusion of single isolated negative links can be interpreted as a continuum percolation problem, where the influence area will depend on the specific underlying topology  [^1] We use the term “topology” in the graph-theoretic sense, referring to connectivity patterns and frustrated cycles in signed graphs, and not in the sense of topological invariants in condensed-matter physics..
 Although this illustration does not resolve our original problem, it offers a valuable physical interpretation of the phenomena that occur as antagonistic interactions are progressively introduced to the system.
@@
 This transition naturally connects to percolation theory, offering a general framework to analyze it.
+The lowest eigenmode thus plays a dual role, capturing both structural fragmentation and the long relaxation bottlenecks induced by near-zero eigenvalues.
 In particular, $|\lambda_{0}\rangle$ captures how the system fragments when global translational invariance –and with it, the unique ultraviolet cutoff, $\Lambda$– is lost.
@@
 This constraint arises from double-precision arithmetic, as discussed in detail in SI.
-Fig.
-LABEL:fig:c:percolation_panel shows the percolation phase transition for a 2D square lattice where a double-peak percolation transition emerges.
+Figure LABEL:fig:c:percolation_panel shows the percolation phase transition for a 2D square lattice where a double-peak percolation transition emerges.
 There, the fraction of points belonging to the largest cluster, $P_{\infty}=n_{\infty}/N$, is the order parameter, with $N$ the total system size and $n_{\infty}$ the number of sites belonging to the percolating cluster.
-The intrinsic fluctuations, $\chi\equiv\sqrt{N}\sigma_{P_{\infty}}$, diverge at $p_{\rm c}$ as expected for usual second-order phase transitions 40.
+The intrinsic fluctuations, $\chi\equiv\sqrt{N}\sigma_{P_{\infty}}$, diverge at $p_{\rm c}$ as expected for usual second-order phase transitions 41.
 As reported in the inset of Fig.
-LABEL:fig:c:percolation_panel, the first peak in the susceptibility, located at $p_{\rm c}\sim 0.10(1)$, is fully compatible with the expected threshold for frustrated plaquettes and/or the emergence of the spin glass phase at zero temperature 41.
-Further examples considering triangular and hexagonal lattices are reported in SI, predicting the threshold where spin glass phases emerge 41, 42 (we will discuss this important point below).
+LABEL:fig:c:percolation_panel, the first peak in the susceptibility, located at $p_{\rm c}\sim 0.10(1)$, is fully compatible with the expected threshold for frustrated plaquettes and/or the emergence of the spin glass phase at zero temperature 42.
+Further examples considering triangular and hexagonal lattices are reported in SI, predicting the threshold where spin glass phases emerge 42, 43 (we will discuss this important point below).
 Moreover, Fig.
-LABEL:fig:e:percolation_panel shows the percolation phase transition for a 3D cubic lattice, where a double peak in the system susceptibility emerges, with $p_{\rm c}=0.22(1)$ 43 (4D analysis are reported in SI).
+LABEL:fig:e:percolation_panel shows the percolation phase transition for a 3D cubic lattice, where a double peak in the system susceptibility emerges, with $p_{\rm c}=0.22(1)$ 44 (4D analyses are reported in SI).
 The cluster-size distribution (see Fig.
@@
 LABEL:fig:f:percolation_panel) confirms that the phase transition effectively belongs to the ordinary percolation universality class, as $P(S)\sim S^{-\tau}$, with $\tau\approx 2.18(1)$ for the 3D case.
-The case of 2D lattices is more subtle, as we observe an exponent $\tau\approx 1.75(1)$, still compatible with previous results in small-to-mid systems in continuum percolation problems, where logarithmic corrections may play a key role in the convergence of the cluster size distribution for low dimensions 44.
+The case of 2D lattices is more subtle, as we observe an exponent $\tau\approx 1.75(1)$, still compatible with previous results in small-to-mid systems in continuum percolation problems, where logarithmic corrections may play a key role in the convergence of the cluster size distribution for low dimensions 45.
 The double peak in the susceptibility represents a delicate point that must be discussed in terms of the geometrical properties of percolation in Euclidean spaces.
@@
 In fact, we observe how the double-peak structure represents a finite-size effect in 2D lattices, which disappears in the infinite-size limit.
-Instead, in 3D, this phenomenon survives in the thermodynamic limit (a fact that is fully compatible with recent observations of the coexistence of spin glass and ferromagnetic order in spin glass 3D systems 46).
+Instead, in 3D, this behavior is consistent with survival in the thermodynamic limit (see SI, and being compatible with observations of coexistence of spin glass and ferromagnetic order in 3D systems 47).
 We report in Fig.
@@
 Note that X defects in the low-density regime are not expected to produce any closed loop, leading to no local frustration of interactions (the system is, then, effectively balanced, i.e., the fundamental eigenvalue is equal to zero).
+Balance, however, does not imply trivial dynamics, as mesoscale fragmentation and competing basins can still emerge.
 This leads to new fundamental dynamic implications that are explored below when X defects are considered together with a dynamic process.
 Figure 3: Ising dynamics.
-(a) Qualitative $T-p$ phase diagram for the 2D squared lattice with S defects.
+(a) Qualitative TpT-p phase diagram for the 2D squared lattice with S defects.
 The lower part shows the energy levels corresponding to the first 100 eigenstates of $\bar{L}$ (see colorbar), for two values of $p$, respectively, in the ferromagnetic phase and the spin-glass one.
@@
 Note that, inside the spin-glass phase, all eigenstates come closer in energy, and their ordering is lost at the ’onset of degeneracy’ at $p_{\rm c}$.
-(b) Qualitative $T-p$ phase diagram for X defects computed with the SL eigenstates through numerical simulations in a 2D square lattice of size $N=64^{2}$.
-The breakup of the ferroelectric phase translates into the emergence of a metastable one where the SL eigenvectors now define multiple degenerate stable states characterized by local order up to the paralectric/ferroelectric phase transition.
+(b) Qualitative TpT-p phase diagram for X defects computed with the SL eigenstates through numerical simulations in a 2D square lattice of size $N=64^{2}$.
+The breakup of the ferroelectric phase translates into the emergence of a metastable one where the SL eigenvectors now define multiple degenerate stable states characterized by local order up to the paraelectric/ferroelectric phase transition.
 At a certain $p$, stable minima becomes unstable, leading to a spin-glass phase.
@@
 Specifically, we have analyzed the dynamical implications of the different topological defects simulating the Edwards-Anderson (EA) Hamiltonian (see Methods),
-$\mathscr{H}=-\sum_{(i,j)}J_{ij}\sigma_{i}\sigma_{j},$ (3)
+$\mathscr{H}=-\underset{\langle i,j\rangle}{\sum}J_{ij}\sigma_{i}\sigma_{j},$ (3)
 where $\sigma_{i}$ is the spin in the $i^{\rm th}$ site, with $J_{ij}=JA_{ij}$, being $A$ the signed adjacency matrix, and $J$ the global coupling strength.
-Fig.
-3(a) and (b) show the temperature phase diagram of the system as a function of the fraction of negative links, $p$, for S and X defects on 2D square lattices, respectively.
-In particular, S defects correspond to the well-known Random Bond Ising Model (RBIM) or ±$\pm J$-model 47, 48 (see Methods).
+Figure 3(a) and (b) show the temperature phase diagram of the system as a function of the fraction of negative links, $p$, for S and X defects on 2D square lattices, respectively.
+In particular, S defects correspond to the well-known Random Bond Ising Model (RBIM) or $\pm J$-model 48, 49 (see Methods).
 Results are achieved through extensive Monte Carlo simulations of spin-spin dynamics, running Eq.
@@
 Note that, unlike the chaotic attractor “surfing” seen in spin glass phases at low temperatures, the MM phase allows access to highly diverse noise-resilient functional configurations.
-Note that, for S defects, the usually reported intriguing re-entrance of the ferroelectric phase to slightly higher values of $p$ when the thermal noise increases 49 can be ascribed to the breaking of small frustrated clusters due to noise effects.
+Note that, for S defects, the usually reported intriguing re-entrance of the ferroelectric phase to slightly higher values of $p$ when the thermal noise increases 50 can be ascribed to the breaking of small frustrated clusters due to noise effects.
 Specifically, thermal noise induces random flips with a single-spin interaction radius that is explicitly smaller than the size of the clustered defects.
 This effect first breaks small spin-clusters, reducing domain wall effects and making more robust the giant ferroelectric cluster.
-Fig.
-3(c) and (d) illustrate the stability of different eigenstates of $\bar{L}$ as the system is heated up to different temperatures.
+Figure 3(c) and (d) illustrate the stability of different eigenstates of $\bar{L}$ as the system is heated up to different temperatures.
 We emphasize that, for X defects, the system can be heated up to the Ising critical temperature, where it is reset by noise, and allow for choosing other possible local minima.
@@
 In the latter case, typical of spin glass behavior, the minima are locally stable, but the system can escape under low noise levels, exploring the phase space (see SI).
-We hypothesize that the spin-glass tree of states 9 is expected to be a superposition of the $N$ eigenvectors of the SL, or pure states, that represent the original set of local minima.
+We hypothesize that the spin-glass tree of states 9 is expected to be a superposition of the $N$ eigenvectors of the SL, or or configurations that can approximate pure states, that represent the original set of local minima.
 It comes natural to introduce topologically-based local order parameters as a function of $T$ that effectively capture the emergent mesoscopic order when spontaneous symmetry breaking forms magnetized spin islands (see Methods).
@@
 3(e) shows the magnetization of the giant cluster of $|\lambda_{2}\rangle$ as initial state, as a function of $p$ and $T$ for the case of X defects.
+We examine a complementary mode to $|\lambda_{0}\rangle$ to probe spectral multiplicity and the presence of competing metastable structures.
 Hence, Fig.
@@
 3(g) shows the narrow temperature range in which the selected starting eigenstates for S defects lose their stability, causing the system to start surfing across other attraction basis, i.e., pure states, or combinations thereof.
-I Topological defects in heterogeneous architectures
+Figure 4: Heterogeneous architectures.
+Order parameter, $P_{\infty}$ versus the fraction of negative links, $p$ for a: (a) ER network with $\langle\kappa\rangle=10$, (b) Small World network from a 2D lattice (rewiring probability $p_{\rm rew}=0.05$) and (c) a diluted 3D lattice ($p_{\rm dil}=0.25$) with normally weighted edges $p\sim\mathcal{N}(1,\sigma)$.
+Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, pp.
+(d) TpT-p Phase diagram of the ER network with a fraction $p$ of X defects measured by the local Ising order parameter $m^{\rm GC}_{\infty}$, and using $\operatorname{sign}(|\lambda_{2}\rangle)$.
+Figure 5: (a) Spectral gap ($\Delta$) and Energy Gap ($E_{0}-E_{1}$) versus $J_{0}/J$ for a fully connected network with symmetric Gaussian interactions, highlighting the closure of the gap at the transition point.
+(b) Activity decay for the antagonistic contact process with a fraction of inhibitory links, $p=0.15$.
+Note the slow decay region indicative of Griffiths phases that emerge beyond $p>0.1$.
+This confirms that the previously identified topological phase transition is consistent with the interpretation that it governs the divergence of dynamical timescales.
+This dynamical slowing down is analytically linked to the proliferation of vanishingly small eigenvalues in the Signed Laplacian spectrum, which emerge as topological defects disrupt the spectral gap.
+Thus, the topological fragmentation of the network provides a structural encoding of the broad distribution of relaxation timescales observed in the dynamics.
+Inset: Temporal decay for a 2D lattice with no inhibitory links ($p=0$), showing standard bistable behavior.
+Parameters: L=96.
+Topological defects in heterogeneous architectures
 To establish whether the SL is able to predict a percolative phase transition in heterogeneous systems, we have analyzed different dilute models and network structures by introducing a random fraction of S defects.
 We highlight that this approach generalizes the detection of frustrated dynamical phases to complex networks and generic disordered systems.
-Specifically, the results for Erdős-Rényi (ER) and Small World (SW) networks are reported in Fig.
+Specifically, the results for ER and Small World (SW) networks are reported in Fig.
 4(a) and (b).
-Additionally, we have analyzed a particularly complex case with significant physical relevance: the 3D diluted cubic lattice, which has originally inspired dilute Ising models 50.
+Additionally, we have analyzed a particularly complex case with significant physical relevance: the 3D diluted cubic lattice, which has originally inspired dilute Ising models 51.
 Now, a fraction $f$ of the existent links are removed, remaining below the edge-percolation threshold to ensure the system is connected.
@@
 Finally, Fig.
-4(d) shows the case of an ER network with X defects reconstructed with the local Ising order parameter.
-Figure 4: Heterogeneous architectures.
-Order parameter, $P_{\infty}$ versus the fraction of negative links, $p$ for a: (a) ER network with $\langle\kappa\rangle=10$, (b) Small World network from a 2D lattice (rewiring probability $p_{\rm rew}=0.05$) and (c) a diluted 3D lattice ($p_{\rm dil}=0.25$) with normally weighted edges $p\sim\mathcal{N}(1,\sigma)$.
-Inset: Variance of the order parameter, $\chi$, as a function of the fraction of negative links, $p$.
-(d) $T-p$ Phase diagram of the ER network with X defects reconstructed with the local Ising order parameter $m^{\rm GC}_{\infty}$ measured on $\operatorname{sign}(|\lambda_{2}\rangle)$.
+4(d) shows the local Ising order parameter for the case of including a fraction $p$ of X defects in an ER network.
+I Implication for E/I models of neural networks
+Antagonistic interactions, central to the balance between excitation and inhibition in neural circuits, have long been a cornerstone of theoretical neuroscience.
+Moving beyond classical spin-lattice approaches that typically give rise to short-range spin glass models, we consider alternative formulations that provide a basis for more refined descriptions of neural computation.
+The first is the Sompolinsky–Crisanti–Sommers (SCS) model 52, which describes the dynamics of large recurrent networks with mixed excitatory and inhibitory synaptic interactions.
+This framework offers a minimal theoretical explanation for the intrinsically irregular, yet collectively organized population dynamics observed in cortical circuits 53.
+The model can be formulated in terms of continuous activity variables $x_{i}(t)$ that interact via quenched couplings $J_{ij}$ and are transformed through a nonlinear transfer function $\phi(\cdot)$; a complete treatment is provided in the Methods.
+Treating the connectivity as quenched disorder and applying dynamical mean-field theory 54, the model predicts a transition from a stable fixed-point regime to high-dimensional chaotic activity when the interaction strength exceeds a critical threshold.
+While the fully asymmetric case yields high-dimensional chaos, the onset of this instability is universally marked by the destabilization of the trivial fixed point (determined by the symmetric part of the interaction spectrum).
+As reported in Fig.
+55(a), we observe that this critical point coincides with the closure of the spectral gap $\Delta$ of the underlying Signed Laplacian.
+Furthermore, the energy gap $E_{0}-E_{1}$ between the ground state and the first excited state of the effective energy landscape vanishes at the same threshold.
+This supports the view that the loss of stability in continuous rate models—regardless of whether the resulting phase is chaotic or glassy—is rooted in the same topological symmetry-breaking mechanism identified in discrete spin systems.
+The second model we consider is a simple two-state (binary) neural network model 55, 56, in which each unit is either active ($s_{i}(t)=1$) or inactive ($s_{i}(t)=0$) at time $t$.
+The state of unit $i$ evolves according to the weighted activity of its $k$ neighbors, $\mathcal{P}_{i}=f\!\left(\Lambda_{i}=\frac{\gamma}{k}\sum_{j}\omega_{ij}s_{j}(t)\right)$, where $\omega_{ij}=\pm 1$ encodes excitatory and inhibitory interactions and $\gamma$ controls the interaction strength.
+To connect with the symmetric interaction framework used in this work, we consider a variant in which excitation and inhibition are assigned to links rather than nodes, so that $\omega_{ij}=\omega_{ji}$ and a fraction $\alpha$ of links are inhibitory.
+This link-based formulation preserves the global E/I balance while ensuring a symmetric coupling matrix, allowing us to isolate the collective effects of antagonistic interactions; full details are provided in the Methods section.
+Fig.
+55(b) shows the temporal decay of the network activity from a fully active state.
+For low fractions of inhibition ($p<0.1$), the system exhibits a single critical point with rapid relaxation or simple bistability (see Inset).
+However, beyond the critical threshold $p\approx 0.1$, we observe the emergence of ’Griffiths phases’–regions characterized by an anomalously slow, power-law decay of activity.
+This phenomenon is not merely a dynamical artifact but a direct consequence of the spectral properties of the Signed Laplacian: as topological defects percolate, they induce a proliferation of degenerate eigenvalues.
+Since relaxation timescales are inversely proportional to these eigenvalues ($\tau\sim 1/\lambda$), the topological fragmentation of the network structurally guarantees the emergence of the diverse, long-lasting timescales often associated with criticality in neural circuits.
 II Discussion and conclusions
-We have solved a decades-long controversy surrounding the nature of spin glasses.
+We provide a new perspective on a decades-long controversy surrounding the nature of spin glasses.
 We have explicitly shown how non-ergodic phases originated by competing interactions, such as the emergence of the spin glass phase at zero temperature, lie in a pure topological symmetry-breaking mechanism that goes beyond the considered dynamics.
@@
 Naturally, such a mechanism strongly depends on the discrete nature of the underlying lattice, leading to different percolation thresholds even if the spatial dimension of the system remains unchanged.
+While the Signed Laplacian does not fully determine nonlinear dynamics, it provides a minimal structural backbone that constrains the landscape on which these dynamics unfold.
 Hence, the SL framework offers a natural approach for detecting spin glass phases in any homogeneous or heterogeneous structure.
@@
 In that case, $d=6$ is the upper dimension for the isotropic percolation universality class, which presents a fractal dimension $d_{f}=4$, which is precisely the upper dimension for the Ising universality class.
-Hence, specific spatial dimensions $d=4$ and $d=5$ are known to exhibit a fractal cluster with $d_{f}\approx 3.04$ and $d_{f}\approx 3.52$, respectively 51, making invalid the mean-field solution for Ising dynamics when competing interactions are considered.
+Hence, specific spatial dimensions $d=4$ and $d=5$ are known to exhibit a fractal cluster with $d_{f}\approx 3.04$ and $d_{f}\approx 3.52$, respectively 57, making invalid the mean-field solution for Ising dynamics when competing interactions are considered.
+Percolation phenomena in correlated and higher-order networked systems have also been investigated in other contexts, such as multiplex and hypergraph models or reaction networks 58, 59.
+While these studies rely on different mechanisms, they are consistent with the broader view that structural organization can influence collective transitions and may motivate future extensions of the present framework.
+One might ask whether other spectral metrics, such as the standard Laplacian or measures of diffusion capacity, could capture these phenomena.
+Standard graph operators, however, typically treat links as purely conductive channels (positive weights), thereby ignoring the crucial algebraic structure of antagonistic interactions.
+While metrics like the Diffusion Capacity describe transport efficiency, they fail to detect the frustration arising from negative cycles.
+The Signed Laplacian is unique in this regard: it naturally incorporates the sign of the interaction into the diffusion process, ensuring that the spectral gap $\Delta$ directly measures the structural balance of the network.
+Consequently, the Signed Laplacian provides a direct and physically interpretable link between the algebraic structure of antagonistic interactions and the effective energy landscape governing relaxation dynamics.
 Finally, our results significantly progress in defining natural stable states, determined by the SL operator, where frustrated systems can exhibit stable minima.
@@
 The discrete signed Laplacian.
-The traditional definition of the Laplacian operator, represented by $L=D-A$, presents serious issues when dealing with signed networks as it is no longer semi-definite positive.
+The traditional definition of the Laplacian operator, represented by L=DAL=D-A, presents serious issues when dealing with signed networks as it is no longer semi-definite positive.
 This is due to the fact that diagonal terms may take null or even negative values and $D^{-1/2}$ and, therefore, $L_{\rm RW}$ is no longer defined.
@@
 In particular, Eq.
-(6) can be reduced, when considering only one negative link, to the problem of point-like particles diffusing in a uniform medium and being absorbed by a single spherical sink 52 (see further details in SI), with $m=m(\bm{x},t)$ the magnetization field,
+(6) can be reduced, when considering only one negative link, to the problem of point-like particles diffusing in a uniform medium and being absorbed by a single spherical sink 60 (see further details in SI), with $m=m(\bm{x},t)$ the magnetization field,
 $\frac{\partial m}{\partial t}=D_{0}\operatorname{\nabla^{2}}m-\lambda(\bm{x})m+s(\bm{x},t),$ (7)
@@
 $G(\textbf{k},\omega)=[-i\omega+D_{0}k^{2}]^{-1}$ (8)
-Through the extension of Eq. (7) to many sinks 52 one can find the following Green function,
+Through the extension of Eq. (7) to many sinks 60 one can find the following Green function,
 $G(\textbf{k},\omega)=[-i\omega+D_{0}k^{2}+nt(\textbf{k},\textbf{k},\omega)]^{-1}$ (9)
@@
 Canonical formulation of signed interactions.
-Once we can consider a semidefinite operator for signed networks, $\bar{L}$, it is possible, in terms of the network propagator, $\hat{K}=e^{-\tau\bar{L}}$, to generalize the ensemble of accessible information diffusion states 31, 33, 32, 53, namely,
+Once we can consider a semidefinite operator for signed networks, $\bar{L}$, it is possible, in terms of the network propagator, $\hat{K}=e^{-\tau\bar{L}}$, to generalize the ensemble of accessible information diffusion states 31, 33, 32, 61, namely,
 $\hat{\rho}(\tau)=\frac{e^{-\tau\bar{L}}}{Z},$ (10)
-where $\rho(\tau)$ is tantamount to the canonical density operator in statistical physics (or to the functional over fields configurations) 35, 54, and $Z(\tau)=\sum_{i}^{N}e^{-\tau\lambda_{i}}$, being $\lambda_{i}$ the set of system’ eigenvalues.
+where $\rho(\tau)$ is tantamount to the canonical density operator in statistical physics (or to the functional over fields configurations) 35, 62, and $Z(\tau)=\sum_{i}^{N}e^{-\tau\lambda_{i}}$, being $\lambda_{i}$ the set of system’s eigenvalues.
 It is possible to, therefore, define the network entropy 31 through the relation
@@
 If $C$ also shows a well-defined plateau, it represents direct evidence of a scale-invariant architecture 33, 36, 38 (such as, for example, 2D lattices).
-In this case, a peak at short times, in turn, independent of system size, captures the ultraviolet cutoff, $\Lambda$ 37.
-This cutoff represents a minimum scale that prevents divergence in usual RG calculations 55 and indicates that the system exhibits some kind of translational invariance if the subsequent plateau is present 38.
+In this case, a peak at short times, in turn, independent of system size, captures the shortest resolvable interaction scale (by analogy with an ultraviolet cutoff), denoted $\Lambda$ in lattice settings 37.
+This cutoff represents a minimum scale that prevents divergence in usual RG calculations 63 and indicates that the system exhibits some kind of translational invariance if the subsequent plateau is present 38.
 Signed Ising dynamics.
-The archetypal spin glass model is defined as the celebrated EA model 48, with the corresponding Hamiltonian
+The archetypal spin glass model is defined as the celebrated EA model 49, with the corresponding Hamiltonian
 $\mathscr{H}=-\sum_{(i,j)}J_{ij}\sigma_{i}\sigma_{j},$ (13)
@@
 Then, the local order parameter runs over the $n_{\infty}$ spins of the giant component of this specific state as indicated by $|\lambda_{i}\rangle$,
-$m^{(i)}_{\infty}(t)=\frac{1}{n^{(i)}_{\infty}}\stackrel{{\scriptstyle[}}{{j}}=1]{n^{(i)}_{\infty}}{\sum}\sigma_{j}(t)$ (17)
+$m^{(i)}_{\infty}(t)=\frac{1}{n^{(i)}_{\infty}}\mathrel{{\mathop{\sum}\limits_{j=1}^{n^{(i)}_{\infty}}}}\sigma_{j}(t)$ (17)
 RBIM vs field theory at $T\to 0$.
 In order to tighten the relation between the fundamental state of the SL and the ground state of the Hamiltonian Eq.
-(3) with the random interaction matrix $J_{ij}$, it is convenient to introduce the field theoretical formulation of the Ising model given by the Hubbard-Stratonovich transformation 56 for which
+(3) with the random interaction matrix $J_{ij}$, it is convenient to introduce the field theoretical formulation of the Ising model given by the Hubbard-Stratonovich transformation 64 for which
 $Z(J)$ $=\sum_{\bm{\sigma}}e^{-\beta\mathscr{H}[\bm{\sigma}]}$ (18)
@@
 $\sum_{i}J_{ki}\left(\phi_{i}-\operatorname{sign}\left[\sum_{l}A_{il}\phi_{l}\right]\right)=0,$
-for $k=1,2,....,N$, where $A_{ij}=J_{ij}/J$ is the adjacency matrix of the network of interactions. The self-consistent solution of the above equation is
+for k=1,2,.,Nk=1,2,....,N, where $A_{ij}=J_{ij}/J$ is the adjacency matrix of the network of interactions. The self-consistent solution of the above equation is
 $\phi_{i}=\operatorname{sign}\left[\sum_{l}A_{il}\phi_{l}\right].$ (19)
@@
 $\sum_{i=1}^{N}A_{ij}\psi_{j}=(k_{i}-\lambda_{0})\psi_{i}\,$, (20)
-where $k_{i}=\sum_{j}|A_{ij}|$ is the coordination number of site $i$ ($k_{i}=2d$ for a $d-$dimensional square lattice). As shown in 57, $\lambda_{0}\leq\min_{i=1}^{N}k_{i}$. This implies that
+where $k_{i}=\sum_{j}|A_{ij}|$ is the coordination number of site $i$ ($k_{i}=2d$ for a $d$-dimensional square lattice). As shown in 65, $\lambda_{0}\leq\min_{i=1}^{N}k_{i}$. This implies that
 $\operatorname{sign}[\psi_{i}]=\operatorname{sign}\left[\sum_{l}A_{il}\psi_{l}\right],$ (21)
@@
 (21) is formally the same of Eq.
-(19) with the fundamental difference that the solution of the former has to be found on the $N-$dimensional spherical surface of radius $\sqrt{N}$ while for the latter has to found on the vertices of the $N-$dimensional cube defined by $\phi_{i}=\pm 1$ for each $i=1,2,...,N$ (note however that in both cases $|\bm{\psi}|=|\bm{\phi}|=\sqrt{N}$.
+(19) with the fundamental difference that the solution of the former has to be found on the NN-dimensional spherical surface of radius $\sqrt{N}$ while for the latter has to found on the vertices of the NN-dimensional cube defined by $\phi_{i}=\pm 1$ for each i=1,2,,Ni=1,2,...,N (note however that in both cases $|\bm{\psi}|=|\bm{\phi}|=\sqrt{N}$.
 This implies that $\operatorname{sign}[\psi_{i}]$ can differ from $\operatorname{sign}[\phi_{i}]$ in some node in particular where there are large fluctuations in the values of $\psi$ between neighboring nodes.
 Anyway, being $\bm{\phi}$ and $\bm{\psi}$ characterized by the same normalization condition, we expect that the binarized vector $\bm{\psi}_{\rm bin}$ of components $\psi_{i}/|\psi_{i}|$ shows a similar behavior with respect the binary vector $\bm{\phi}$ on a suitable coarse-grained scale as large as the frustration walls.
+Sompolinsky–Crisanti–Sommers (SCS) rate model
+We consider a recurrent network of $N$ units with continuous activity variables $x_{i}(t)$, generated from effective inputs $h_{i}(t)$ through a nonlinear transfer function,
+$x_{i}(t)=\phi\!\left(g\,h_{i}(t)\right),$ (22)
+where $g$ is a gain parameter and $\phi(x)=\tanh(x)$ unless stated otherwise. The inputs evolve according to first-order relaxation dynamics,
+$\frac{\mathrm{d}h_{i}(t)}{\mathrm{d}t}=-h_{i}(t)+\sum_{j=1}^{N}J_{ij}\,x_{j}(t),$ (23)
+where $J_{ij}$ are quenched synaptic couplings drawn as
+$J_{ij}\sim\mathcal{N}\!\left(0,\,\frac{g^{2}}{N}\right),\qquad J_{ii}=0.$ (24)
+In the limit $N\to\infty$, the recurrent input to each neuron becomes an effective Gaussian process, and dynamical mean-field theory (DMFT) yields a self-consistent equation for the autocorrelation $C_{x}(t,t^{\prime})=\langle x(t)x(t^{\prime})\rangle$.
+This theory predicts a transition at $g_{c}=1$: for $g<g_{c}$ the network converges to a stable fixed point, while for $g>g_{c}$ the dynamics is chaotic.
+The structure of the connectivity matrix determines whether the dynamics is relaxational.
+For generic (asymmetric) $J$,  Eq.
+(23) does not derive from an energy function, and the dynamics may exhibit chaotic behavior.
+However, when the couplings are symmetric, $J_{ij}=J_{ji}$, the dynamics becomes a gradient descent in an energy landscape,
+$\frac{\mathrm{d}h_{i}(t)}{\mathrm{d}t}=-\frac{\partial\mathcal{E}}{\partial h_{i}}\Big|_{h(t)},$ (25)
+with energy
+$\mathcal{E}(h_{1},\dots,h_{N})$ $=\sum_{i=1}^{N}\int_{0}^{h_{i}}\mathrm{d}h\,\phi(gh)$ (26)
+$\quad-\frac{1}{2}\sum_{i,j=1}^{N}J_{ij}\,\phi(gh_{i})\,\phi(gh_{j}).$
+The first term encodes the single-unit activation nonlinearity, while the second term represents pairwise synaptic interactions.
+Binary excitatory–inhibitory model with symmetric interactions
+We also consider a binary-state network model in which each node is either active ($s_{i}(t)=1$) or inactive ($s_{i}(t)=0$) at discrete time $t$.
+Each node interacts with its $k$ neighbors through weighted inputs.
+The probability of activation is given by
+$\mathcal{P}\big(s_{i}(t{+}1)=1\big)=f\!\left(\Lambda_{i}(t)\right),\qquad\Lambda_{i}(t)=\frac{\gamma}{k}\sum_{j}\omega_{ij}s_{j}(t),$
+where $\gamma$ controls the global coupling strength and $\omega_{ij}\in\{\pm 1\}$ encodes excitatory (+1+1) or inhibitory (1-1) interactions. The activation function is taken to be piecewise linear:
+$\begin{cases}0,&\Lambda<0,\\ \Lambda,&0\leq\Lambda\leq 1,\\[3.0pt] 1,&\Lambda>1.\end{cases}$
+To connect directly with the symmetric interaction framework used in the present work, we consider a link-symmetrized variant in which inhibition and excitation are assigned to links rather than nodes.
+The network is taken to be random-regular and undirected, so that each node has exactly $k$ neighbors and $\omega_{ij}=\omega_{ji}\in\{\pm 1\}$.
+A fraction $p$ of links are inhibitory and (1p)(1{-}p) are excitatory, ensuring that the coupling matrix $(\omega_{ij})$ is symmetric by construction.
+The update rule remains unchanged, but inhibition is now encoded at the level of pairwise interactions rather than neuronal types, providing a minimal symmetric framework for a contact–process–like antagonistic coupling.